OpenStudy (anonymous):

Find the slope of the line passing through the points (-2, 9) and (5, 3). A. -6/7 B. 4 C. -7/6 D. 1/4

OpenStudy (anonymous):

@luckythebest

OpenStudy (anonymous):

If (-2,9) = (x1,y1) and (5,3) = (x2,y2) Slope = y2-y1/x2-x1.

OpenStudy (anonymous):

i dont understand how to do this?????

OpenStudy (anonymous):

Ok. You need to know the concept of SLOPE. If a point in the co-ordinate plane shifts from one position to another, then the difference in the y-co-ordinate divided by the difference in x-coordinate is called SLOPE. So if (-2,9) moves to (5,3) , Then the SLOPE = 9-3/-2-5 (in the same direction. i.e. DONT DO 9-3/5+2; which is wrong)

OpenStudy (anonymous):

so then the answer is C -7/6

OpenStudy (anonymous):

@ilovedaivd haha nope. am pretty sure that 9-3 = 6 and -2-5 = -7 making it -6/7 :p

OpenStudy (anonymous):

that what i thought firts but how does the 6 become a negative?

OpenStudy (anonymous):

first*

OpenStudy (anonymous):

6/-7 , or -6/7, mean the same. You're just shifting the negative sign above which is perfectly fine.

OpenStudy (anonymous):

oo okay cause i got 6/-7 but i didnt see it in my answers X) thank you